OpenStudy (anonymous):

solve using differentiation: a triangle has a base 12 feet long and an altitude 8 feet high. find the area of the largest rectangle that can be inscribed in the triangle so that the base of the rectangle falls on the base of the triangle. answer is 24 sq. ft. *needs solution

OpenStudy (goformit100):

@uri

OpenStudy (anonymous):

is it a right triangle? an equilateral triangle, or just some triangle?

OpenStudy (anonymous):

given the base and the altitude as 12 ft and 8 feet respectively,

OpenStudy (anonymous):



OpenStudy (anonymous):

so that the base of the rectangle falls to the base of triangle. How would be the figure of that..

OpenStudy (anonymous):

the point \(y\) labelled on the right satisfies the equation \(y=-3x+12\) the length of the base of the rectangle is \(2x\) and so the area of the rectangle is \[A(x)=2x(-3x+12)=-6x^2+24x\]

OpenStudy (anonymous):

you don't need calculus to find the max of that quadratic find the first coordinate of the vertex that will give you \(x\)

OpenStudy (anonymous):

so how can we come up with 24 sq ft? sorry i'm really not good at this. :(

OpenStudy (anonymous):

first of all is it clear how i came up with \(y=-3x+12\) ?

OpenStudy (anonymous):

Because of the equation of the side?

OpenStudy (anonymous):

to answer your question, how do you come up with \(24\) the vertex of the quadratic, \[A(x)=-6x^2+24x\] is \((2,24)\) and there is the 24 if you want to use calculus to find the vertex, take the derivative of \(A(x)=-6x^2+24\) and get \(A'(x)=-12x+24\) set it equal to zero and solve, and get \(x=2\) then you find that \(A(2)=24\)

OpenStudy (anonymous):

yes, the equation of the line through \((0,12)\) and \((4, 0)\) is \(y=-3x+12\)

OpenStudy (anonymous):

the real key here is to stick the triangle inside a coordinate axis in order to get an equation for the area without that, it is almost impossible once you use a coordinate axis, the equation for the area is not that hard to find

OpenStudy (anonymous):

hm.. i think i get it already. wait i'll just going to solve for it on my own. thankyou for the help.

OpenStudy (anonymous):

can i ask you another problem later?

OpenStudy (anonymous):

yw sure you can ask later

OpenStudy (anonymous):

Thank you. :)

OpenStudy (anonymous):

can you show me the solution to this y=−3x+12.? i think i'm getting it wrong -_-

OpenStudy (anonymous):

you can do that in your head the line from \((0,12)\) to \((4,0)\) has \(y\) intercept 12 right?

OpenStudy (anonymous):

as for the slope, you go over 4, down 12, so slope is \(-\frac{12}{4}=-3\)

OpenStudy (anonymous):

like with slope \(-3\) and \(y\) intercept \(12\) is \(y=-3x+12\)

OpenStudy (anonymous):

i have another solution ..\[y=-\frac{ 12 }{ 8 }x +12\] would give me a x=4 and then substituting it to our equation of the line give me y=6 Area of rec = xy = (4)(6)=24